Stephanie Corporation sells a single product. Budgeted sales for the year are anticipated to be 662,000...

  1. Stephanie Corporation sells a single product. Budgeted sales for the year are anticipated to be 662,000 units, estimated beginning inventory is 103,000 units, and desired ending inventory is 88,000 units. The quantities of direct materials expected to be used for each unit of finished product are given below.

    Material A 0.50 lb. per unit @ $0.53 per pound
    Material B 1.00 lb. per unit @ $2.31 per pound
    Material C 1.20 lb. per unit @ $0.79 per pound

    The dollar amount of material A used in production during the year is

    a. $171,455

    b. $613,356

    c. $1,494,570

    d. $175,430
Answer #1

Answer :A) $ 171,455

Estimated Production in Units = Sales in Units + Ending Inventory -Beginning Inventory

Units Prodcued = 662,000+88,000-103,000 = 647,000

Materila A Used in Production = 647,000*0.53 *0.50= $171,455 (Answer)